OpenStudy (anonymous):

An electric fan is set on its HIGH setting. After the Low push button is depressed, the angular speed of the fan blades decreases to a value of 8.00x10^2 rev/min in 1.75sec. The deceleration is 42.0 rad/s^2. Determine the initial angular speed of the blades in rev/min

OpenStudy (mos1635):

a=Δω/Δt a=2π*Δf/Δt a=2π*(f2-f1)/Δt ..... f1=f2-a*Δt/2π f2=800/60 rev/sec a=-42 rad/sec^2 Δt=1.75sec so f1=25 rev/sec f1=1500 rev/min